Kyd and North are playing a game. Kyd selects one card from a standard 52-card deck. If Kyd selects a face card (Jack, Queen, or King), North pays him $5. If Kyd selects any other type of card, he pays North $2.

a) What is Kyd's expected value for this game? Round your answer to the nearest cent. $

b) What is North's expected value for this game? Round your answer to the nearest cent. $

c) Who has the advantage in this game?
1
Expert's answer
2020-06-30T08:22:39-0400

The probability of selecting face card (jack,queen or king)= (4+4+4)/52 = 12/52 = 0.2308

probability of not selecting a face is 1-0.2308 = 0.7692

a)Kyd's Expected value = 0.2308*5 - 0.7692*2 = -0.3844

B)North's expected value is = 0.7692*2 -0.2304*5 =0.3844

c) North has advantage since he has a higher expected value.